# Hawaii: Hierarchical Visual Knowledge Transfer for Efficient Vision-Language Models

### v1

@@ Abstract
Improving the visual understanding ability of vision-language models (VLMs) is crucial for enhancing their performance across various tasks. While using multiple pretrained visual experts has shown great promise, it often incurs significant computational costs during training and inference. To address this challenge, we propose Hawaii, a novel framework that distills knowledge from multiple visual experts into a single vision encoder, enabling it to inherit the complementary strengths of several experts with minimal computational overhead. To mitigate conflicts among different teachers and switch between different teacher-specific knowledge, instead of using a fixed set of adapters for multiple teachers, we propose to use teacher-specific Low-Rank Adaptation (LoRA) adapters with a corresponding router. Each adapter is aligned with a specific teacher, avoiding noisy guidance during distillation. To enable efficient knowledge distillation, we propose fine-grained and coarse-grained distillation. At the fine-grained level, token importance scores are employed to emphasize the most informative tokens from each teacher adaptively. At the coarse-grained level, we summarize the knowledge from multiple teachers and transfer it to the student using a set of general-knowledge LoRA adapters with a router. Extensive experiments on various vision-language tasks demonstrate the superiority of Hawaii, compared to the popular open-source VLMs.
@@ 1 Introduction
Fine-grained distillation. As different teachers’ knowledge differ due to the heterogeneity of their training data, architecture, and optimization methods, in MoLA, teacher-specific LoRA adapters are employed to avoid the conflicts between teachers’ knowledge. Each adapter is aligned with each teacher separately, allowing the student encoder to learn from diverse teachers while mitigating noisy distillation. Moreover, to emphasize the informative tokens generated by each teacher, at the fine-grained level, HKD utilizes a new token importance scoring method, which assigns weights to tokens as indicated by visual tokens and text tokens.
Coarse-grained distillation. To obtain the collective consensus among visual teachers, HKD summarize the knowledge from multiple teachers using a projector. Then, MoLA incorporates a set of general-knowledge LoRA adapters and a router to align the student with the collective consensus for a global alignment.
@@
  • Extensive experiments on various vision-language tasks (gurari_vizwiz_2018, ; singh_towards_2019, ; hudson_gqa_2019, ; lu2022learn, ; li_evaluating_2023, ; fu_mme_2024, ; leonardis_mmbench_2025, ; yue_mmmu_2024, ; kembhavi_diagram_2016, ; li_seed-bench_2024, ) show that Hawaii achieves better performance on all the benchmark datasets compared to the baseline model (LLaVA-1.5 (liu_improved_2023, )). Specifically, the performance on VizWiz, SQA, and MMBench is improved by 3.9, 3.7, and 2.6, respectively.
@@ 2 Hawaii
[…]
[width=1]pics/model.pdf
Figure 1: The overall architecture of Hawaii. We use two teachers for simplicity. (a) MoLA (Section 2.2) consists of teacher-specific LoRA adapters (Teacher Adp.) and general-knowledge LoRA adapters (General Adp.) with two routers controlling the activation of adapters. (b) Coarse-grained distillation (Section 2.3.1) first summarizes the knowledge from multiple teachers and then transfers it to the student encoder globally. “T1 Feat.”, “T2 Feat,”, and “Sum. T. Feat.” represents the visual features $I^{T}_{*}$ generated by different teachers and the summarized teacher features $I^{T}_{\textit{cg}}$. (c) In the fine-grained distillation (Section 2.3.2), teacher-specific LoRA adapters (T. Adp.) and token importance scoring (Figure 2) are employed to select and learn from the most informative tokens.
In this section, we introduce our new framework, i.e., Hawaii, which learns from multiple powerful visual teachers for a better visual perception ability. Hawaii inherit the complementary strengths of several experts without incurring substantial computational overhead. First, we introduce the architecture of Hawaii. Second, we present the details of MoLA, which consists of a set of teacher-specific and general-knowledge LoRA adapters in Section 2.2. Last, we provide the details of our hierarchical knowledge distillation method, which contains the coarse- and fine-grained distillation in Section 2.3.
@@ 2.1 Architecture
The overall architecture of Hawaii is presented in the top part of Figure 1. It follows the general design (vision expert-projector-LLM) of existing MLLMs (chen_internvl_2024, ; liu_visual_2023, ; shieagle, ). The vision expert is trained to distill knowledge from multiple pretrained vision experts and produces visual tokens used for visual comprehension. The projector maps the visual tokens to the LLM input space, and the LLM generates the instruction-following response.
The vision encoder $f_{v}(\cdot)$ takes the input image and generates a set of visual tokens $I^{S}\in\mathcal{R}^{m\times D}$, where $m$ is the number of visual tokens and $D$ is the dimension of each token. To boost the performance, instead of using multiple vision encoders (wang2025leo, ; shieagle, ; azadani2025leo, ), which would be computationally expensive, only one student vision encoder is employed. And, it is trained to distill knowledge from multiple pretrained vision experts (kirillov_segment_2023, ; oquab2023dinov2, ). We introduce the mixture-of-LoRA-adapter (MoLA, Section 2.2) module that enables the student encoder to learn from diverse teachers in a fine-grained (Section 2.3.2) and coarse-grained (Section 2.3.1) manner.
@@
An LLM $f_{\textsc{LLM}}(\cdot)$ then takes the mapped visual tokens $f_{p}(I^{S})$ and the textual instruction tokens $T$ as input to generate the instruction-following response $Y=\{y_{i}\}_{i\in[L]}$ as,
@@
where $L$ is the length of the response, and $y_{<i}$ is the previous tokens of $y_{i}$.
@@ 2.2 Mixture of LoRA Adapters
Directly fine-tuning the student encoder is challenging, as it often leads to overfitting on the limited fine-tuning data and catastrophic forgetting (movekd2025, ; jia_visual_2022, ). To this end, we propose a mixture-of-LoRA-adapter (MoLA) module consisting of teacher-specific LoRA adapters and general-knowledge LoRA adapters (edward_j_hu_lora_2021, ) to enable the student encoder to learn from diverse teachers without forgetting, as illustrated in Figure 1 (a).
Teacher-specific LoRA adapters. Learning from multiple teachers is challenging (luo_knowledge_2019, ; shen_customizing_2019, ; shen_amalgamating_2019, ), as each teacher (kirillov_segment_2023, ; oquab2023dinov2, ; fang_eva_2023, ) might have different training data, model architecture, and training objectives. Directly transferring diverse teachers’ knowledge to the student could lead to noisy distillation and performance drop. To this end, we introduce a set of teacher-specific LoRA adapters $\{a_{i}^{T}\}_{i\in[N_{t}]}$. Each adapter is designed to align with one teacher only, which avoids the conflicts between multiple teachers (see Section 2.3.2). Those adapters are applied to each feedforward layer of the student encoder $f_{v}(\cdot)$, where $N_{t}$ is the number of teachers.
General-knowledge LoRA adapters. For learning the collective consensus from teachers and the training data, we introduce a set of general-knowledge LoRA adapters $\{a_{i}^{G}\}_{i=1}^{N_{g}}$ that are applied to each feedforward layer of the student encoder $f_{v}(\cdot)$, where $N_{g}$ is the number of general-knowledge LoRA adapters. The details of this general (global) knowledge transfer are provided in Section 2.3.1.
We adopt the general (sparse) design of mixture-of-experts (MoE) (wu_mixture_2023, ; dai_deepseekmoe_2024, ) to select the LoRA adapters based on the hidden inputs of each layer. Specifically, we employ two sparse routers, i.e., $f_{r}^{T}(\cdot)$ and $f_{r}^{G}(\cdot)$, to select the teacher-specific LoRA adapters and general-knowledge LoRA adapters, respectively. Formally, for each feedforward layer of the student encoder, the MoE output $F^{*}(\cdot)$ is computed as,
@@
$w.\quad i=\mathop{\textsc{argmax}}(f_{r}^{T}(h)),j=\mathop{\textsc{argmax}}(f_{r}^{G}(h))\,,$
where $h$ is the hidden input of the current layer, $F(\cdot)$ is the current layer, and $\mathop{\textsc{argmax}}(\cdot)$ is the argmax operation. We denote the visual tokens generated by the student encoder with MoLA as $I^{S}$.
@@ 2.3 Hierarchical Knowledge Distillation
To integrate diverse teachers’ knowledge into a single student encoder, we propose a hierarchical knowledge Distillation (HKD) mechanism that transfers knowledge at two levels of granularity, i.e., coarse-grained and fine-grained levels. Specifically, for coarse-grained distillation (Section 2.3.1), we summarize the knowledge from multiple teachers (collective consensus) and transfer it to the student encoder globally. For fine-grained distillation (Section 2.3.2), teacher-specific LoRA adapters are employed to align with each teacher separately for a precise noise transfer. Moreover, to attend to the most informative tokens during knowledge transfer, we introduce a token importance scoring method (Figure 2) based on the similarity among teachers’ visual tokens and the input instructions.
@@ 2.3.1 Coarse-Grained Distillation (CGKD)
To obtain the collective consensus, i.e., summarized teacher feature, each teacher’s visual features are first unshuffled (chen_internvl_2024, ; shieagle, ; shi_real-time_2016, ) to have the same length (chen_internvl_2024, ) as the student’s visual features $I^{S}\in\mathcal{R}^{m\times D}$. Then, those visual tokens are channel-wise concatenated and the summarized feature $I^{T}_{cg}$ is obtained by applying a two-layer MLP $f_{cg}(\cdot)$ as,
@@
Next, we apply the coarse-grained distillation loss $\mathcal{L}_{\text{cg}}$ to transfer the collective consensus by minimizing the mean square error (MSE) between the summarized features $I^{T}_{cg}$ and the student encoder output $I^{S}$, as,
$\mathcal{L}_{\text{cg}}=\mathop{\textsc{MSE}}(I^{S},I^{T}_{cg})\,,$ (4)
[…]
@@ 2.3.2 Fine-Grained Distillation (FGKD)
Using LoRA adapters (edward_j_hu_lora_2021, ) to transfer knowledge from one teacher to a student has been proven to be successful. However, transferring knowledge from multiple teachers to a single student is challenging, especially when using a fixed set of LoRA adapters (movekd2025, ) for all the teachers. The reason is that the noisy and redundant teachers’ knowledge can hinder the learning process and lead to suboptimal performance (luo_knowledge_2019, ; shen_customizing_2019, ; shen_amalgamating_2019, ), due to the conflicts among teachers, which arises from the heterogeneity of training data, architectures, and the training algorithms.
To address this challenge, we propose the fine-grained distillation (FGKD) that leverages teacher-specific LoRA adapters and token importance scoring. Each teacher-specific LoRA adapter is designed to align with one teacher only, allowing the student to learn from each teacher separately. Token importance scoring is used to select and attend to the most informative tokens from each teacher during knowledge transfer, reducing the noise and redundancy.
@@
[…]
[width=]pics/score.pdf
Figure 2: The calculation of token importance score $s_{i}$. To focus on the most informative tokens, we consider the similarity among teacher’s features and the input instructions $T$.
Token importance scoring. The key to knowledge distillation is to transfer the most important information (hinton2015distilling, ). As previous studies show that not all tokens are equally informative (shang_llava-prumerge_2024, ; wang2025leo, ; zhang_llava-mini_2025, ; movekd2025, ), to identify the most informative tokens, we introduce a new similarity-based importance score that considers teachers’ visual tokens and the input instructions $T$, allowing us to prioritize tokens that are more relevant to the task context. Specifically, for the $i$-th teacher, we compute the token importance score $s_{i}\in\mathcal{R}^{1\times m}$ as,
@@
where $\hat{I}^{T}_{i}\in\mathcal{R}^{m\times D}$ and $\hat{T}$ are the visual tokens and input instructions projected by a learnable two-layer MLP to have the same dimension of the student features, $D$ is the dimension of the visual tokens, $\mathop{\textsc{Concat}}(\cdot)$ is the concatenation operation, $\mathop{\textsc{Mean}}(\cdot)$ is the mean operation, and $\mathop{\textsc{Softmax}}(\cdot)$ is the softmax operation.
Now, with the token importance scores $\{s_{i}\}_{i\in[m]}$, the fine-grained distillation loss $\mathcal{L}_{\text{fg}}$ is calculated as,
@@ 2.4 Training Objectives
The overall training objective of Hawaii is to minimize the loss consisting of the text generation loss $\mathcal{L}_{\text{gen}}$ (touvron_llama_2023, ; touvron_llama2_2023, ), the coarse-grained distillation loss $\mathcal{L}_{\text{cg}}$ (Equation 4), the fine-grained distillation loss $\mathcal{L}_{\text{fg}}$ (Equation 6), and the MoE balance loss $\mathcal{L}_{\text{mb}}$ (dai_deepseekmoe_2024, ; liu_routers_2024, ; dou_loramoe_2024, ) as,

### v2

@@ Abstract
Improving the visual understanding ability of vision-language models (VLMs) is crucial for enhancing their performance across various tasks. While using multiple pretrained visual experts has shown great promise, it often incurs significant computational costs during training and inference. To address this challenge, we propose Hawaii, a novel framework that distills knowledge from multiple visual experts into a single vision encoder, enabling it to inherit the complementary strengths of several experts with minimal computational overhead. To mitigate conflicts among different teachers and switch between different teacher-specific knowledge, instead of using a fixed set of adapters for multiple teachers, we propose to use teacher-specific Low-Rank Adaptation (LoRA) adapters with a corresponding router. Each adapter is aligned with a specific teacher, avoiding noisy guidance during distillation. To enable efficient knowledge distillation, we propose fine-grained and coarse-grained distillation. At the fine-grained level, token importance scores are employed to emphasize the most informative tokens from each teacher adaptively. At the coarse-grained level, we summarize the knowledge from multiple teachers and transfer it to the student using a set of general-knowledge LoRA adapters with a router. Extensive experiments on various vision-language tasks demonstrate the superiority of Hawaii compared to popular open-source VLMs. The code is available at https://github.com/yimuwangcs/wise-hawaii.
@@ 1 Introduction
Fine-grained distillation. As each teacher’s knowledge is different, due to the heterogeneity of training data, architecture, and optimization methods, in MoLA, teacher-specific LoRA adapters are employed to avoid conflicts between teachers’ knowledge. Each adapter is aligned with its teacher separately, allowing the student encoder to learn from diverse teachers while mitigating noisy distillation. Moreover, to emphasize the informative tokens generated by each teacher, at the fine-grained level, HKD utilizes a new token importance scoring method, which assigns weights to tokens according to the similarity to the text instructions and visual features.
Coarse-grained distillation. To obtain the collective consensus among visual teachers, HKD summarizes the knowledge from multiple teachers using a projector. Then, MoLA incorporates a set of general-knowledge LoRA adapters and a router to align the student with the collective consensus for a global alignment.
@@
  • Extensive experiments on various vision-language tasks (gurari_vizwiz_2018, ; singh_towards_2019, ; hudson_gqa_2019, ; lu2022learn, ; li_evaluating_2023, ; fu_mme_2024, ; leonardis_mmbench_2025, ; yue_mmmu_2024, ; kembhavi_diagram_2016, ; li_seed-bench_2024, ) show that Hawaii achieves better performance on all the benchmark datasets compared to the baseline model (LLaVA-1.5 (liu_improved_2023, )). In particular, the performance on VizWiz, SQA, and MMBench is improved by 7.8%, 5.5%, and 4.0%, respectively.
@@ 2 Hawaii
Figure 1: The overall architecture of Hawaii. We use two teachers for simplicity. (a) MoLA (Section˜2.2) consists of teacher-specific LoRA adapters (Teacher Adp.) and general-knowledge LoRA adapters (General Adp.) with two routers controlling the activation of adapters. (b) Coarse-grained distillation (Section˜2.3.1) first summarizes the knowledge from multiple teachers and then transfers it to the student encoder globally. “T1 Feat.”, “T2 Feat,”, and “Sum. T. Feat.” represents the visual features $I^{T}_{*}$ generated by different teachers and the summarized teacher features $I^{T}_{\textit{cg}}$. (c) In the fine-grained distillation (Section˜2.3.2), teacher-specific LoRA adapters (T. Adp.) and token importance scoring (Figure˜2) are employed to select and learn from the most informative tokens.
In this section, we introduce the Hawaii framework, which learns from multiple powerful visual teachers for a better visual perception ability. Hawaii inherits the complementary strengths of several experts without incurring substantial computational overhead. First, we introduce the architecture of Hawaii. Second, we present the details of MoLA, which consists of a set of teacher-specific and general-knowledge LoRA adapters in Section˜2.2. Last, we provide the details of our hierarchical knowledge distillation method, which contains the coarse- and fine-grained distillation in Section˜2.3.
@@ 2.1 Architecture
The overall architecture of Hawaii is presented in the upper part of Figure˜1. It follows the general design (vision expert-projector-LLM) of existing MLLMs (chen_internvl_2024, ; liu_visual_2023, ; shieagle, ). The vision expert is trained to distill knowledge from multiple pretrained vision experts and produces visual tokens used for visual comprehension. The projector maps the visual tokens to the LLM input space, and the LLM generates the instruction-following response.
The vision encoder $f_{v}(\cdot)$ takes the input image and generates a set of visual tokens $I^{S}\in\mathcal{R}^{m\times D}$, where $m$ is the number of visual tokens and $D$ is the dimension of each token. To boost the performance, instead of using multiple vision encoders (wang2025leo, ; shieagle, ; azadani2025leo, ), which would be computationally expensive, only one student vision encoder is employed. And, it is trained to distill knowledge from multiple pretrained vision experts (kirillov_segment_2023, ; oquab2023dinov2, ). We introduce the mixture-of-LoRA-adapter (MoLA, Section˜2.2) module that enables the student encoder to learn from diverse teachers in a fine-grained (Section˜2.3.2) and coarse-grained (Section˜2.3.1) manner.
@@
An LLM $f_{\textsc{LLM}}(\cdot)$ then takes the mapped visual tokens $f_{p}(I^{S})$ and the textual instruction tokens $T$ as input to generate the instruction-following response $Y=\{y_{i}\}_{i\in[L]}$ as
@@
where $L$ is the length of the response and $y_{<i}$ is the previous tokens of $y_{i}$.
@@ 2.2 Mixture of LoRA Adapters
Directly fine-tuning the student encoder is challenging, as it often leads to overfitting on the limited fine-tuning data and catastrophic forgetting (movekd2025, ; jia_visual_2022, ). To avoid this, we propose a mixture-of-LoRA-adapter (MoLA) module consisting of teacher-specific LoRA adapters and general-knowledge LoRA adapters (edward_j_hu_lora_2021, ) to enable the student encoder to learn from diverse teachers without forgetting. MoLA is illustrated in Figure˜1 (a).
Teacher-specific LoRA adapters. Learning from multiple teachers is challenging (luo_knowledge_2019, ; shen_customizing_2019, ; shen_amalgamating_2019, ), as each teacher (kirillov_segment_2023, ; oquab2023dinov2, ; fang_eva_2023, ) might have different training data, model architecture, and training objectives. Directly transferring diverse teachers’ knowledge to the student could lead to noisy distillation and performance drop. To avoid this, we introduce a set of teacher-specific LoRA adapters $\{a_{i}^{T}\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of teachers. Each adapter is designed to align with one teacher only, which avoids the conflicts between multiple teachers (see Section˜2.3.2). Those adapters are applied to each feedforward layer of the student encoder $f_{v}(\cdot)$.
General-knowledge LoRA adapters. For learning the collective consensus from teachers and the training data, we introduce a set of general-knowledge LoRA adapters $\{a_{i}^{G}\}_{i=1}^{N_{g}}$ that are applied to each feedforward layer of the student encoder $f_{v}(\cdot)$, where $N_{g}$ is the number of general-knowledge LoRA adapters. The details of this general (global) knowledge transfer are provided in Section˜2.3.1.
We adopt the general (sparse) design of mixture-of-experts (MoE) (wu_mixture_2023, ; dai_deepseekmoe_2024, ) to select the LoRA adapters based on the hidden inputs of each layer. Specifically, we employ two sparse routers, i.e., $f_{r}^{T}(\cdot)$ and $f_{r}^{G}(\cdot)$, to select the teacher-specific LoRA adapters and general-knowledge LoRA adapters, respectively. Formally, for each feedforward layer of the student encoder, the MoE output $F^{*}(\cdot)$ is computed as
@@
$\text{with~}i=\mathop{\textsc{argmax}}(f_{r}^{T}(h))\text{~and~}j=\mathop{\textsc{argmax}}(f_{r}^{G}(h)),$
where $h$ is the hidden input of the current layer and $F(\cdot)$ is the current layer. We denote the visual tokens generated by the student encoder with MoLA as $I^{S}$.
@@ 2.3 Hierarchical Knowledge Distillation
To integrate diverse teachers’ knowledge into a single student encoder, we propose a hierarchical knowledge Distillation (HKD) mechanism that transfers knowledge at two levels of granularity, i.e., coarse-grained and fine-grained levels. Specifically, for coarse-grained distillation (Section˜2.3.1), we summarize the knowledge from multiple teachers (collective consensus) and transfer it to the student encoder globally. For fine-grained distillation (Section˜2.3.2), teacher-specific LoRA adapters are employed to align with each teacher separately for a precise noise transfer. Moreover, to attend to the most informative tokens during knowledge transfer, we introduce a token importance scoring method (Figure˜2) based on the similarity among teachers’ visual tokens and the input instructions.
@@ 2.3.1 Coarse-Grained Distillation (CGKD)
To obtain the collective consensus, i.e., summarized teacher feature, each teacher’s visual features are first unshuffled (chen_internvl_2024, ; shieagle, ; shi_real-time_2016, ) to have the same length (chen_internvl_2024, ) as the student’s visual features $I^{S}\in\mathcal{R}^{m\times D}$. Then, those visual tokens are channel-wise concatenated and the summarized feature $I^{T}_{cg}$ is obtained by applying a two-layer MLP $f_{cg}(\cdot)$ as
@@
Next, we apply the coarse-grained distillation loss $\mathcal{L}_{\text{cg}}$ to transfer the collective consensus by minimizing the mean square error loss (MSE) between the summarized features $I^{T}_{cg}$ and the student encoder output $I^{S}$ as
$\mathcal{L}_{\text{cg}}=\mathop{\textsc{MSE}}(I^{S},I^{T}_{cg}).$ (4)
@@ 2.3.2 Fine-Grained Distillation (FGKD)
Using LoRA adapters (edward_j_hu_lora_2021, ) to transfer knowledge from one teacher to a student has proven to be successful. However, transferring knowledge from multiple teachers to a single student is challenging, especially when using a fixed set of LoRA adapters (movekd2025, ) for all the teachers. The reason is that the noisy and redundant teachers’ knowledge can hinder the learning process and lead to suboptimal performance (luo_knowledge_2019, ; shen_customizing_2019, ; shen_amalgamating_2019, ), due to the conflicts among teachers, which arises from the heterogeneity of training data, architectures, and the training algorithms.
To address this challenge, we propose the fine-grained distillation (FGKD) that exploits teacher-specific LoRA adapters and token importance scoring. Each teacher-specific LoRA adapter is designed to align with one teacher only, allowing the student to learn from each teacher separately. Token importance scoring is used to select and attend to the most informative tokens from each teacher during knowledge transfer, reducing the noise and redundancy.
@@
Figure 2: The calculation of token importance score $s_{i}$. To focus on the most informative tokens, we consider the similarity among the teacher’s features and the input instructions $T$.
Token importance scoring. The key to knowledge distillation is to transfer the most important information (hinton2015distilling, ). As previous studies show that not all tokens are equally informative (shang_llava-prumerge_2024, ; wang2025leo, ; zhang_llava-mini_2025, ; movekd2025, ), to identify the most informative tokens, we introduce a new similarity-based importance score that considers teachers’ visual tokens and the input instructions $T$, allowing us to prioritize tokens that are more relevant to the task context. Specifically, for the $i$-th teacher, we compute the token importance score $s_{i}\in\mathcal{R}^{1\times m}$ as
@@
where $\hat{I}^{T}_{i}\in\mathcal{R}^{m\times D}$ and $\hat{T}$ are the visual tokens and input instructions projected by a learnable two-layer MLP to have the same dimension of the student features. $D$ is the dimension of the visual tokens.
Now, with the token importance scores $\{s_{i}\}_{i\in[m]}$, the fine-grained distillation loss $\mathcal{L}_{\text{fg}}$ is calculated as
@@ 2.4 Training Objectives
The overall training objective of Hawaii is to minimize the loss consisting of the text generation loss $\mathcal{L}_{\text{gen}}$ (touvron_llama_2023, ; touvron_llama2_2023, ), the coarse-grained distillation loss $\mathcal{L}_{\text{cg}}$ (Equation˜4), the fine-grained distillation loss $\mathcal{L}_{\text{fg}}$ (Equation˜6), and the MoE balance loss $\mathcal{L}_{\text{mb}}$ (dai_deepseekmoe_2024, ; liu_routers_2024, ; dou_loramoe_2024, ). This is given as